Blood pressure is a measure of the blood force against the walls of the arteries.
Blood pressure that is consistently more than 140/90 mm Hg is considered high,
but for people with diabetes, 130/80 mm Hg is considered high. Normal blood
pressure is below 120/80 mm Hg. The larger value represents the pressure when
the heart contracts and pushes blood out (systolic), and the smaller value is the
lowest pressure when the heart relaxes between beats (diastolic). Consider a
population of a non-diabetic males with a mean systolic blood, in which 19%
are classified as having high blood pressure. Assume that the systolic blood
pressure is normally distributed.
1. Compute the standard deviation of the systolic blood pressure.
2. Compute the 26-th and 76-th percentiles of the systolic blood pressure.
3. Stage 1 hypertension is defined as a systolic blood pressure between 140
mm Hg and 160 mm Hg, or a diastolic blood pressure between 90 mm
Hg and 100 mm Hg. Among 100 randomly selected individuals from this
population, what is the probability that 95% of them are classified as
having stage 1 hypertension according to the value of the systolic blood
pressure, ignoring the value of the diastolic blood pressure?
4. Stage 2 hypertension is defined as a systolic blood pressure larger than
160 mm Hg, or a diastolic blood pressure larger than 100 mm Hg. If four
individuals are randomly selected from this population, what is the probability
that their underlying sample mean is classified as having stage 2
hypertension according to the value of the systolic blood pressure, ignoring
the value of the diastolic blood pressure.
Answer #1

ANSWER: Given that 94 115 Signal= 28.4 Q-95.978 pg134.028 have systolic blood pressuse move than 40 194. 19) X=115 +7C sigma)X = 115 H40.67) (28.4) Q1 = 95.972 Qys 75th percentile 759, values are less than Q3 1= x-01 sigma 0.67 - X-11928.4 X=115+0.6